OpenStudy (anonymous):

@iGreen 7. x/4=12/6 a. x=6 b. x=12 c. x=8 d. x=48 8. 9=m-4 a. m=36 b. m=13 c. m=5 d. m=9/4 9. j-3=-7 a. j=-10 b. j=21 c. j=-4 d. j=4

OpenStudy (igreen):

x/4 = 12/6 Multiply 4 to both sides: 12/6 * 4 = 8 x = 8

OpenStudy (igreen):

9 = m - 4 Add 4 to both sides: 4 + 9 = 13 m = 13

OpenStudy (igreen):

j - 3 = -7 Add 3 to both sides: -7 + 3 = -4 j = -4

OpenStudy (anonymous):

10. you need to make 18 cupcakes for a birthday party. you have already made 11. how many mpre cupcakes do you need to make? write and equation and solve

OpenStudy (igreen):

You need 18 cupcakes. You have already made 11. x + 11 = 18 is your equation. To solve it, subtract 11 to both sides. 18 - 11 = 7 x = 7 So you need to make 7 more cupcakes.